### v1

# supplement

Abstract
Launching and focusing phonon-polaritons present novel opportunities for low-loss guiding of subdiffractionally confined light at the nanoscale. Despite significant efforts to improve control in polaritonic media, focused and spatially confined phonon-polariton waves have only been achieved in the in-plane anisotropic crystals (such as MoO3) and remain elusive in the in-plane isotropic crystals (such as hexagonal boron nitride). In this study, we present a previously unexplored approach to launching phonon-polaritons by employing hBN subwavelength resonators coupled to gold nanoantennas. The integration of gold nanoantennas enables efficient coupling to the far-field component of mid-infrared light, while the geometry of the resonators defines the wavefront curvature, spatial confinement, and focusing of the launched polaritons. Using standard lithographic protocols, we achieve strong field enhancement and resonant mode localization, generating phonon-polaritons with precise spatial definition. Scattering-type scanning near-field optical microscopy reveals the real-space optical contrast of these modes. This innovative and practical approach introduces a new paradigm for fabricating nanoresonators that actively launch and focus phonon-polaritons, opening avenues for advanced nanophotonic devices.
phonon-polariton, polaritonic optics, strong confinement, hexagonal boron nitride, focusing
Introduction. Strong light-matter interactions play a crucial role in well-established photonic devices, such as lasers, integrated photonic circuits, sensors, and hold promise for advancing future photonic quantum technologies, entangled spin-photon interfaces, quantum repeaters, and quantum memories1. In particular, interactions between photons and phonons, which are quantized vibrations of a crystal lattice, result in the formation of propagating and localized phonon-polaritons (PhPs). PhPs have garnered significant attention as they confine light substantially below the diffraction limit2, 3, 4. PhP resonances occur in various crystalline materials within their Reststrahlen band — a region of the spectrum between transverse (TO) and longitudinal (LO) optical phonons, where the materials exhibit properties of a hyperbolic medium5. This enables PhP waveguides and cavities with exceptional electromagnetic confinement in the mid-infrared range (mid-IR) and with remarkably low propagation loss compared to plasmon-polaritons 6. The mid-IR range of the spectrum is highly sought after due to its potential applications in medical diagnostics 7, 8, bio- and gas sensing 9, 10, 11, LiDaRs for spacecraft 12, 13, 14, optomechanics 15, optical trapping 16, and other fields. To achieve micrometer-scale integration of devices in the mid-IR range, the strong confinement provided by hyperbolic materials is likely the only viable solution.
Figure 1: Diversity of PhP waves. a) Schematics of the setup for the s-SNOM measurements of PhPs in hBN sub-wavelength resonators. Note that the far-field light covers the majority of the sample’s area (spot radius $\geq$ 10 $\mathrm{\mu}$m) b) Optical amplitude (S4) image of the 45-nm-thick hBN flake placed on the pre-patterned Si/SiO2 substrate with Au stripes. Colored dotted lines indicate line-cuts of various propogating PhP waves. c) S4 profiles taken across corresponding propagating PhP waves. d) S-SNOM image of the cavity and propagating PhPs at various excitation wavelengths. e) S4 profiles taken across the Au stripe and the cavity at $\omega$ = 1410 cm-1.
Hexagonal boron nitride (hBN) is one of the most promising and technologically accessible materials that provides a hyperbolic optical response in the mid-IR region of the spectrum. hBN is a uniaxial hyperbolic material that exhibits two types of hyperbolic optically active transitions — Type I - 760-820 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}>0;\varepsilon_{zz}<0$) and Type II - 1360-1600 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}<0;\varepsilon_{zz}>0$) 3, 17, where $\varepsilon_{ij}$ is the permittivity tensor components. Recent studies have reported on the reliable growth of large-scale high-quality crystals with the possibility to control their isotopic purity 18, 19, 20. Currently, plasmonic nanoantennas are widely used to couple the far-field component of light with polaritonic systems, facilitating the selective launch of phonon-polaritons (PhPs)21, 22, 23, 24. However, this approach suffers from a critical limitation: the lack of control over the wavefront and spatial confinement of the launched PhPs. While hBN nanocavities exhibit remarkable mid-IR light confinement (down to $\lambda$/500, where $\lambda$ is the incident wavelength), an extraordinary Purcell factor (up to 1012), and a high quality factor ($Q>400$), the resonant modes in these cavities remain trapped within the resonators25, 26, 27, 28. Up to our best knowledge, there has been no demonstration of utilizing these localized modes as sources for focused PhPs. This poses a significant promise to the advancement of PhP-based technologies. In particular, effective control over spatially confined propagation and focusing in uniaxial hyperbolic materials ($\varepsilon_{xx}=\varepsilon_{yy}$), such as hBN, remains elusive. Although spatially confined PhP focusing along the optical axis has been achieved in biaxial hyperbolic materials like MoO3 29, This method is fundamentally constrained by crystallographic directions, making it unsuitable for arbitrary lensing applications.
Here, we demonstrate hBN sub-wavelength resonators coupled to gold nanoantennas, compatible with standard lithographic protocols and capable of launching focused and spatially confined long-lived PhPs. We fabricate sub-wavelength hBN cavities at the edges between Au and Si++/SiO2 (285 nm) substrates. The Au edge serves as a nanoantenna, enabling coupling to the far-field component of light, while the Au/SiO2 interface acts as a partially transparent wall for PhPs: on the SiO2 side, it localizes a resonant mode within the cavity, while on the Au side, it promotes the launch of resonantly enhanced PhPs into the hBN bulk. The localized modes exhibit pronounced field enhancement and generate polaritons spatially defined by the cavity dimensions. Using scattering-type scanning near-field optical microscopy (s-SNOM), we visualize the real-space optical contrast and distinguish between cavity-launched ($\lambda_{p}$) and tip-launched ($\lambda_{p}$/2) modes. The former results from coupling with the far-field component of mid-IR light, which makes a significant contribution to the PhPs launched in this system. The far-field component, resonantly enhanced in the nano-cavities, acquires sufficient momentum to launch polaritons into the bulk of the hBN/Au heterostructures via scattering processes at the SiO2/Au interface. This approach contrasts with previous demonstrations of focused PhPs, where polaritons in hBN were primarily excited by the s-SNOM tip 30. The shapes of our resonators determine the focal points and wavefronts of the PhPs. By fabricating nano-resonators in various shapes, we demonstrate "concave lens"-like and "convex lens"-like focusing. "Concave lens"-like cavities achieve confinement down to $\lambda/55$ at the focal point. Thus, we propose a novel method for launching and focusing spatially confined in-plane PhPs, fully compatible with standard lithographic procedures.
Results. In this work, we study two types of samples: (i) mechanically cleaved hBN crystals (of 30-45 nm thickness) transferred on the pre-patterned Au on Si++/SiO2 (285 nm) substrates as is (see Methods) and (ii) nanoresonators fabricated at the edges between SiO2 and Au using electron-beam lithography and plasma etching (see the schematics in Fig.1). Fig. 1a shows the measurement scheme for s-SNOM accompanied by the schematics of the fabricated sample. Here, we use a commercially available s-SNOM equipped with PtIr-coated (23 nm thick) AFM tips with the curvature radii less than 25 nm. A tunable ($\lambda$ = 5.5-11.5 $\mu$m) mid-infrared quantum cascade laser allows us to cover Type II hyperbolic transition - 1360-1600 cm-1 ($\epsilon_{xx}=\epsilon_{yy}<0;\epsilon_{zz}>0$) (see Methods section for details).
Figure 2: Focusing of PhPs using patterned nanoresonantors. a) AFM topography of the Au edge with the resonators of triangular shapes. The hBN flake thickness is 45 nm. Red dashed line shows the edge between Au and SiO2. b) Experimental s-SNOM amplitude (S3) map vs. the tip position recorded at $\omega$ = 1450 cm-1. c) s-SNOM amplitude (S3) profiles taken across cavity-lanched waves (red, blue, and black lines) and the hBN flake edge on Au (green line). Dashed lines show exponential decay fit for each curve.
Fig. 1b shows optical amplitude (s4) versus tip position on a 42-nm-thick hBN flake placed on top of a 50-nm-thick Au stripe at $\omega$ = 1452 cm-1 (see topography and optical phase data in Extended Data Fig. S1). We focus on three specific regions: i) to the left of the Au stripe, ii) on top of the Au stripe, and iii) to the right of the Au stripe. The first region comprises a triangular hBN piece, where we observe distinct maxima and minima of the optical amplitude (related to the electric field amplitude $(\textrm{Re E}_{\textrm{z}})^{2}$), characteristic of localized resonant PhP modes28. In the second region, we identify two types of the PhP waves. The first type appears near the edges of the hBN flake (highlighted by the orange dotted line). The second wave emanates from the cavity (red dotted line). In the third region, additional two PhP waves are visible. The first wave is positioned near the edges of the hBN flake (blue dotted line). The second wave extends from the edge of the Au stripe to the right across the entire hBN bulk (green dotted line).
The waves, which are localized close to the edge and propagate perpendicularly to it, are the tip-launched PhPs (we do not observe hBN edge-launched fringes in our geometry)31. These waves originate from the mid-IR light scattered on the AFM tip. The observed fringes represent only a half-wavelength of the corresponding PhP ($\lambda_{PhP}$/2) since the detection moves along with the excitation 32, 33. Based on this, we attribute the orange and the blue PhP waves to the tip-launched modes on Au and SiO2, respectively.
Fig. 1c shows linecuts taken along the dotted lines in the panel Fig. 1b. As expected, the red and the green curves exhibit doubled wavelengths $\lambda_{PhP}$ (compared to the tip-launched waves on Au and SiO2, respectively). Intriguingly, they also demonstrate notably longer propagation lengths. We attribute these PhPs to the nanocavities-launched waves 33. These waves originate from the light scattered at the edge of the Au stripe. In the case of the green curves, the infrared light launches the corresponding PhP in the hBN on SiO2. For the red curves, we assume that the Au edge launches the PhPs inside the cavity where they localize and enhance, and then it leaks back through the semi-transparent interface into the hBN bulk on Au. We emphasize that the tip-launched PhPs (orange and blue curves) only exist when the AFM tip is present in the measurement setup, while the PhPs, indicated by the red and green lines, couple to the far-field light and exist as long as there is far-field light excitation.
Fig.1d shows a set of zoomed-in maps of the cavity-launched wave at various excitation wavelengths. The launched PhP (white arrow) moves and shrinks along with the resonant maximum position inside the cavity (see the detailed investigation of the dispersion in Extended Data Fig. S2). Fig.1e shows optical amplitude profiles taken along the red and the orange lines in Fig.1d at $\omega$ = 1410 cm-1. We find that the electric field is significantly enhanced inside the cavity exhibiting a two-fold larger reflectivity compared to the bare gold signal (and seven-fold larger compared to hBN/SiO2 signal). This observation establishes our nanocavities as effective PhP resonators.
To investigate PhP waves’ dependence on their shapes and sizes, we utilize a standard electron beam lithography and plasma etching (see Methods). We fabricate nanocavities of the desired shapes at the edges of gold patterns. Fig. 2a shows a topography of one of our samples. Attributed to an exceptional alignment (within $\sim$10 nm) between the e-beam lithography steps, we create the matching edges between the patterned hBN and the Au substrates. This allows us to avoid unwanted leakages of the polaritonic modes away from our nanocavities. On the left hand side of the orange dashed line, a 45 nm thick hBN flake covers a 50 nm thick Au stripe. On the right hand side, the patterned hBN triangles hang from the Au edge and lie down on the SiO2 substrate (note a smooth hBN topography height transition across the Au/SiO2 edge). Respectively, Fig. 2b shows an experimental s-SNOM amplitude (S3) taken at $\omega$ = 1450 cm-1 (see its dispersion and the analogous set of data for a 31 nm thick hBN in Extended Data Fig. S3 and S4, respectively). Inside the cavities, the scattered signal maxima confirm the field localization and enhancement.
Figure 3: Numerical simulations of PhPs alongside comparisons to experimental data. a) Numerically calculated dispersion of propagating PhP modes generated at the hBN/SiO2 and hBN/Au interfaces (solid lines) compared to the experiment (symbols). Insets show the experimental s-SNOM optical amplitude (S4) maps and the theoretical spatial distributions of $(\textrm{Re E}_{\textrm{z}})^{2}$ in triangular cavities, for the wavelength indicated by the grey dashed lines. b) Numerically and experimentally obtained dispersion of the attenuation constant (Im k) for hBN/SiO2 and hBN/Au interfaces. The error bars indicate the spread of Im k in the experimental data. c) $(\textrm{Re E}_{\textrm{z}})^{2}$ in triangular cavities which we experimentally probe in Fig. 2. The distributions are taken at the cavity’s resonant frequencies. All calculations are done for hBN thickness of 45 nm.
Figure 4: Concave- and convex-type PhPs focusing. a, b, c, d) AFM topography, numerical modeling, s-SNOM optical amplitude (S3) maps and the schematics of "concave lens"-like cavity. e, f, g, h) Same data for the "convex lens"-like cavity. Light yellow dashed lines show the edge between Au and SiO2. Experimental s-SNOM optical amplitude (S3) maps recorded at $\omega$ = 1410 cm-1. The thickness of hBN is 55 nm. Red dashed lines are a guide for eyes indicating boundaries of cavity-launched polaritons. Dotted lines highlight maxima of optical amplitude for cavity-launched and tip-launched PhPs. Green lines demonstrate tip-launched PhP waves.
Similar to the case in Fig. 1, two different PhP waves can be observed. The shorter one is a tip-launched PhP (indicated by a green arrow with the wavelength $\lambda_{PhP}/2$). This wave results from mid-infrared light scattered by the AFM tip, which excites phonon-polaritons that propagate toward the edge of the sample, reflect back, and interfere with themselves32, 34, 35, 36, 37. Interestingly, we observe the second PhP wave originating from the cavities themselves. Here, the observe that the peak-to-peak distance corresponds to the whole $\lambda_{PhP}$. It is worth noting that $\lambda_{PhP}/2$ fringes are completely suppressed in front of the cavities, and only $\lambda_{PhP}$ fringes are observed.
Importantly, we find that the cavity-launched PhPs acquire well-defined focal points. For the top cavity (cavity 1), the PhP shrinks before and expands after the focal point (marked by the red arrow). The lateral size of the focal point corresponds to a confinement of $\lambda/25$. For the bottom two cavities, PhPs shrink with the distance away from the cavities towards their focal points that are located outside of the detectable regions. The orange dashed lines serve as guides to the eye, indicating the boundaries of the cavity-launched phonon-polaritons. At the focal point, we identify a "hot spot" that corresponds to the electric field node. Consequently, the PhP’s width and the focal point location can be varied by the size of a cavity.
Fig. 2c shows linecuts taken across the dashed lines in Fig. 2b (solid lines) and their exponential decay fits (dashed lines). Tip-launched PhPs show a strong first peak of oscillation but a fast decay of the PhP amplitude due to high losses 38. On the contrary, cavity-launched PhPs exhibit a stable amplitude of oscillations with the considerably slow decay. As a result, pronounced optical amplitude oscillations of the cavity-launched PhPs are propagating more than three times farther than the tip-launched PhPs.
[…]
Numerical simulations. To shed light on the physics of PhPs in our cavities and to establish an approach for their engineering with the desirable characteristics, we build a theoretical model that predicts optical properties of PhPs and effectively replicates the PhP dynamics in the triangular cavities observed in the experiment. Fig. 3a shows the dispersion of the propagation constant (Re k) of the PhP guided modes at the hBN/SiO2 and hBN/Au interfaces. The thickness of the hBN flake is 45 nm. The numerically calculated dispersion appears in a good agreement with the data extracted from the experiment albeit the dispersion for the hBN/Au interface appears to be slightly off. We attribute it to the buckling of the hBN flakes on the Au edge and to the surface roughness of Au; both can create gaps between hBN and Au that reduce the effective refractive index at the interface.
Fig. 3b shows the dispersion of attenuation constant Im k and compares it to the experimental data on the decay rate $\kappa=2\times\textrm{Im k}$. The numerical and experimental attenuation constants well-agree with each other between 1400 cm-1 and 1460 cm-1 but diverge at the frequencies above 1460 cm-1. We attribute this divergence to the absorption on PMMA (e-beam lithography resist) residues that remain on the sample surface after the fabrication process. In this spectral range, PMMA exhibits stronger absorption with an increase in wavenumber39, 40. See more data on the dispersion, refractive index and the attenuation constant for the 31- and 42-nm thick hBN in Extended Data Fig. S5 and S6. We use the complex effective refractive index extracted from the numerical data in Fig. 3a,b to construct a 2D model. This model captures the PhP wave dynamics in the cavities and waveguides formed at the hBN/SiO2 and hBN/Au interfaces, correspondingly. As can be seen from the insets in Fig. 3b, the simulated field distributions are in good agreement with the experimental optical amplitude maps. However, in the experiment, the optical amplitude along the gold edge is noticeably higher, which can also be attributed to the buckling effect that cannot be accounted for in our 2D model. Fig. 3c demonstrates the spatial distributions of the PhPs localized in the triangular hBN/SiO2 cavities and propagating in the hBN/Au waveguide. The distributions show that the triangular cavities can support localized PhP modes that "leak" into the hBN/Au environment and that the leaked PhP waves are focused at a distance from the triangle base depending on the triangle size, which is also observed in the experiment (see Fig. 2b). This observation indicates that the optical phenomena in phonon-polaritons excited within a 3D structure can be effectively described by a 2D model.
The schematic analysis of the triangular and rectangular cavities (see Extended Data Fig. S7) shows that the shape of the PhP’s wavefront is closely related to the geometry of the cavity. For example, the size of the triangular cavity determines the focal point, while cavity-launched PhPs of rectangular cavities do not shrink and, as a result, do not obtain a focal point. Conclusively, due to a strong confinement, polaritons obey the laws of the wave optics, which brings up and opportunity to realize "concave lens"-like and "convex lens"-like focusing.
PhP lensing. Fig. 4a, 4b, and 4c show the AFM topography, numerically calculated electric field $(\textrm{Re E}_{\textrm{z}})^{2}$, and the experimental near-field optical amplitude (S3) of a submicron "concave lens"-like cavity, respectively. This patterned shape exhibits defocusing of the polaritons. We provide an intuitive explanation for this process in Fig. 4d. The PhP waves launched by the Au edge, propagate perpendicular to the Au edge inside the cavity and reflect back from the concave cavity edge. Upon reaching the cavity’s edge, they reflect normally to the curvature of the reflecting plane. The negative curvature of the reflecting plane results in the negative curvature of the polariton wavefront and, thus, defocusing.
Figs. 4e-g show a cavity that operates in a "convex lens" regime, focusing polaritons at a point distant from the cavity, similar to the triangular cavities. One can observe a focal point where the curvature of the wavefront changes sign from positive to negative as it passes through the focal point. Fig. 4h illustrates a schematics of the "convex lens"-like focusing. The position of the focal point and the confinement may be adjusted by the size of the cavity. For a 0.6-$\mu$m-radius cavity, the confinement reaches $\lambda/55$ (see Extended Data Fig. S8). Furthermore, several cavities launching criss-crossing PhPs can be used for the formation of the interference patterns (see Extended Data Fig. S9) forming a lattice of $(\textrm{Re E}_{\textrm{z}})^{2}$ minima and maxima with a period equal to the polariton wavelength at the intersection. Thus, using the interference of cavity-launched polaritons with various wavefronts, it is possible to form a variety of polaritonic lattices for structured light-matter interaction and opto-mechanical purposes.
Discussion. To accelerate the technological application of spatially defined PhP modes, it is necessary to find a way to precisely control their spatial distribution and focus them at specific real-space locations. Some substantial progress has been made in tuning and focusing of the PhPs using in-plane anisotropic hyperbolic materials such as MoO341, 42, 43. In the case of the uniaxial hyberbolic hBN crystals, most efforts have involved a use of substrates with an in-plane anisotropy in the relevant spectral ranges 44, 45, 46. However, the influence of these substrates is generally limited. An elegant solution has been proposed by Chaudhary et. al. 30 by using phase-fluid materials (Ge3Sb2Te6 on a CaF2 substrate) that locally change their dielectric constants and thus facilitate polaritonic waveguides and metalenses. However, the focusing effect remained rather hindered and the scalable application is somewhat elusive.
In conclusion, we have developed hBN sub-wavelength resonators that are compatible with standard lithographic techniques, enabling the launch of focused, long-lived PhPs. We fabricate the subwavelength hBN cavities on the interfaces between Au and SiO2 partially transparent for PhPs. The localized modes showed a significant increase of the electric field and launched polaritons that remain confined in the real-space defined by the shapes and sizes of the cavities. Using s-SNOM, we studied the real-space optical contrast of the cavity-launched PhPs. Based on the PhP’s wavelength, we demonstrated coupling to the far-field component of the mid-IR light. The far-field component, which is resonantly enhanced in the cavities, gains enough momentum to launch polaritons into the bulk of the hBN/Au heterostructures through the scattering process at the SiO2/Au interfaces. The shapes of our resonators determine the focal point location as well as the wavefront shapes. By creating different forms of nano-resonators, we showed a focusing effect similar to a "concave lens" and a "convex lens." In particular, the "concave lens"-like cavities enable field confinement down to $\lambda/55$ at the focal points. By this, we realize a novel technique for generating and focusing spatially confined easy-plane PhPs compatible with the standard silicon technology.
Methods
Device fabrication.
Si++/SiO2 substrates were pre-patterned using standard photolithography and metal deposition procedures. For photolithography, SUSS MicroTec MJB4 mask aligner and SPR-700 photoresist were used. After the photoresist development, we deposited metal (Ti/Au, 10/50 nm) using the Oerlikon 450B electron beam vacuum deposition system.
hBN flakes were prepared via mechanical exfoliation using PDMS adhesive tapes and were transferred on Si/SiO2 substrates for selection. Relying on optical contrast, we selected flakes of suitable thicknesses and transferred them on the pre-patterned Si++/SiO2 substrates using the PMDS/PC dry transfer method47, 48.
For the cavity fabrication, we used the Raith EBPG5200 electron beam lithographer and PMMA 950 A4 e-beam resist. After the resist development, exposed hBN was etched out using Oxford ICP-RIE plasma etcher with CHF3 plasma source (40 sscm, 90 mTorr, 60 W).
Near-field optical measurements.
We used a commercially available scattering-type scanning near-field microscope (s-SNOM) developed by Neaspec/Attocube (cryo-neaSCOPE) with a pseudo-heterodyne mode to a detect a background-free optical response ($\Omega\pm NM$) using the decoupled optical amplitude and phase49, 50, 51. PtIr-coated AFM tips (Nanoworld, 23 nm coating) with a resonant frequency $\approx$ 250 kHz were used to obtain the topography and the optical maps. AFM tips were illuminated by a tunable commercially available mid-infrared quantum cascade laser (QCL) using 1-2 mW of power (MIRcat, $\lambda$ = 5.5-11.5 $\mu$m from the DRS Daylight Solutions Inc.) All optical measurements presented in the manuscript were obtained on the 3rd or the 4th harmonics (i.e., S3 and S4, respectively) to ensure the background-free detection.
Numerical model
To describe the optical phenomena in PhPs at the hBN/SiO2 and hBN/Au interfaces, we employ numerical calculations in COMSOL Multiphysics, Wave Optics module. The dispersions of the propagation and attenuation constants of the PhP modes alongside the complex effective refractive indices of these modes are calculated in a 3-layer structure, where the top layer is presented by air, the second layer is presented by hBN which is an anisotropic (along the hard axis) material with permittivity tensor taken from 52 and optical axis directed perpendicular to the layers, and the third layer is either SiO2 or Au with refractive indices taken from 53, 54. The obtained effective refractive indices are used as refractive indices for hBN/SiO2 and hBN/Au domains in the 2D model, and the surrounding space is treated as a dielectric with a refractive index of 1. As a source for the 2D model, we specify the surface current density at the edges of the hBN/SiO2, hBN/Au domains, and the surrounding space.
Acknowledgements
We thank Andrey Bogdanov, Lorenzo Orsini, Frank H. L. Koppens and Hanan Herzig Sheinfux for helpful discussions. The authors thank Morten R. Eskildsen for his assistance in setting up the experiment. P.S. acknowledges support from the start-up fund provided by the University of Notre Dame (#373837). K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Author contributions
P.S. and B.B conceived and designed the experiment. B.B. fabricated samples and performed the measurements. The data were analyzed and interpreted by B.B., S.L. and P.S. The hBN crystals were provided by K.W. and T.T. S.L performed the numerical calculations. B.B, S.L., and P.S. wrote the manuscript. P.S. supervised the work.
Competing Financial Interests
The authors declare no competing financial interests.
Data Availability Statement
The data that support the findings of this study is available from the corresponding authors upon reasonable request.

### v2

@@
 Abstract
-Launching and focusing phonon-polaritons present novel opportunities for low-loss guiding of subdiffractionally confined light at the nanoscale.
-Despite significant efforts to improve control in polaritonic media, focused and spatially confined phonon-polariton waves have only been achieved in the in-plane anisotropic crystals (such as MoO3) and remain elusive in the in-plane isotropic crystals (such as hexagonal boron nitride).
-In this study, we present a previously unexplored approach to launching phonon-polaritons by employing hBN subwavelength resonators coupled to gold nanoantennas.
-The integration of gold nanoantennas enables efficient coupling to the far-field component of mid-infrared light, while the geometry of the resonators defines the wavefront curvature, spatial confinement, and focusing of the launched polaritons.
-Using standard lithographic protocols, we achieve strong field enhancement and resonant mode localization, generating phonon-polaritons with precise spatial definition.
-Scattering-type scanning near-field optical microscopy reveals the real-space optical contrast of these modes.
-This innovative and practical approach introduces a new paradigm for fabricating nanoresonators that actively launch and focus phonon-polaritons, opening avenues for advanced nanophotonic devices.
+Phonon-polaritons offer significant opportunities for low-loss, subdiffractional light guiding at the nanoscale.
+Despite extensive efforts to enhance control in polaritonic media, focused and spatially confined phonon-polariton waves have only been realized in in-plane-anisotropic crystals (e.g., MoO3) and remain elusive in in-plane-isotropic materials (e.g., hexagonal boron nitride, hBN).
+In this study, we introduce a novel approach to launching phonon-polaritons by leveraging hBN subwavelength cavities at the Au/SiO2 interface, enabling efficient coupling of cavities to the far-field component of mid-infrared light.
+Utilizing standard lithographic techniques, we fabricated subwavelength cavities of various shapes and sizes, demonstrating strong field enhancement, resonant mode localization, and generation of propagating phonon-polaritons with well-defined spatial structure.
+The cavity geometry governs wavefront curvature, spatial confinement, and polariton focusing, providing control over their propagation and achieving record-high in-plane confinement up to $\lambda/70$.
+Scattering-type scanning near-field optical microscopy reveals the real-space optical contrast of these cavity-launched modes, allowing for detailed characterization.
+We believe that our cavity-based approach to phonon-polariton focusing in isotropic media will pave the way for advanced nanophotonic applications.
 phonon-polariton, polaritonic optics, strong confinement, hexagonal boron nitride, focusing
 Introduction.
-Strong light-matter interactions play a crucial role in well-established photonic devices, such as lasers, integrated photonic circuits, sensors, and hold promise for advancing future photonic quantum technologies, entangled spin-photon interfaces, quantum repeaters, and quantum memories1.
-In particular, interactions between photons and phonons, which are quantized vibrations of a crystal lattice, result in the formation of propagating and localized phonon-polaritons (PhPs).
-PhPs have garnered significant attention as they confine light substantially below the diffraction limit2, 3, 4.
-PhP resonances occur in various crystalline materials within their Reststrahlen band — a region of the spectrum between transverse (TO) and longitudinal (LO) optical phonons, where the materials exhibit properties of a hyperbolic medium5.
+Strong light-matter interactions play a crucial role in well-established photonic devices, such as lasers, integrated photonic circuits, sensors, and hold promise for advancing future photonic quantum technologies, single-photon sources, entangled spin-photon interfaces, and quantum memories1.
+In particular, strong light-matter interactions can be achieved in propagating and localized phonon-polaritons (PhPs), which result from coupling between photons and phonons, quantized vibrations of a crystal lattice.
+PhP resonances occur in various crystalline materials within Reststrahlen band – a region of the spectrum between transverse (TO) and longitudinal (LO) optical phonons, where the materials exhibit properties of a hyperbolic medium2.
+PhPs have garnered significant attention as they confine light substantially below the diffraction limit3, 4, 5.
 This enables PhP waveguides and cavities with exceptional electromagnetic confinement in the mid-infrared range (mid-IR) and with remarkably low propagation loss compared to plasmon-polaritons 6.
-The mid-IR range of the spectrum is highly sought after due to its potential applications in medical diagnostics 7, 8, bio- and gas sensing 9, 10, 11, LiDaRs for spacecraft 12, 13, 14, optomechanics 15, optical trapping 16, and other fields.
+The mid-IR range of the spectrum is highly sought after due to its potential applications in medical diagnostics 7, 8, bio- and gas sensing 9, 10, 11, LiDaRs for spacecraft 12, 13, 14, optomechanics 15, and optical trapping 16.
 To achieve micrometer-scale integration of devices in the mid-IR range, the strong confinement provided by hyperbolic materials is likely the only viable solution.
@@
 a) Schematics of the setup for the s-SNOM measurements of PhPs in hBN sub-wavelength resonators.
-Note that the far-field light covers the majority of the sample’s area (spot radius $\geq$ 10 $\mathrm{\mu}$m) b) Optical amplitude (S4) image of the 45-nm-thick hBN flake placed on the pre-patterned Si/SiO2 substrate with Au stripes.
+Note that the far-field light covers the majority of the sample’s area (spot radius $\geq$ 10 $\mathrm{\mu}$m) b) Optical amplitude (S4) image of the 42-nm-thick hBN flake placed on the pre-patterned Si/SiO2 substrate with Au stripes.
 Colored dotted lines indicate line-cuts of various propogating PhP waves.
@@
 Hexagonal boron nitride (hBN) is one of the most promising and technologically accessible materials that provides a hyperbolic optical response in the mid-IR region of the spectrum.
-hBN is a uniaxial hyperbolic material that exhibits two types of hyperbolic optically active transitions — Type I - 760-820 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}>0;\varepsilon_{zz}<0$) and Type II - 1360-1600 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}<0;\varepsilon_{zz}>0$) 3, 17, where $\varepsilon_{ij}$ is the permittivity tensor components.
+hBN is a uniaxial hyperbolic material that exhibits two types of hyperbolic optically active transitions — Type I - 760-820 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}>0;\varepsilon_{zz}<0$) and Type II - 1360-1600 cm-1 ($\varepsilon_{xx}=\varepsilon_{yy}<0;\varepsilon_{zz}>0$) 4, 17, where $\varepsilon_{ij}$ is the permittivity tensor components.
 Recent studies have reported on the reliable growth of large-scale high-quality crystals with the possibility to control their isotopic purity 18, 19, 20.
-Currently, plasmonic nanoantennas are widely used to couple the far-field component of light with polaritonic systems, facilitating the selective launch of phonon-polaritons (PhPs)21, 22, 23, 24.
-However, this approach suffers from a critical limitation: the lack of control over the wavefront and spatial confinement of the launched PhPs.
-While hBN nanocavities exhibit remarkable mid-IR light confinement (down to $\lambda$/500, where $\lambda$ is the incident wavelength), an extraordinary Purcell factor (up to 1012), and a high quality factor ($Q>400$), the resonant modes in these cavities remain trapped within the resonators25, 26, 27, 28.
-Up to our best knowledge, there has been no demonstration of utilizing these localized modes as sources for focused PhPs.
-This poses a significant promise to the advancement of PhP-based technologies.
-In particular, effective control over spatially confined propagation and focusing in uniaxial hyperbolic materials ($\varepsilon_{xx}=\varepsilon_{yy}$), such as hBN, remains elusive.
-Although spatially confined PhP focusing along the optical axis has been achieved in biaxial hyperbolic materials like MoO3 29, This method is fundamentally constrained by crystallographic directions, making it unsuitable for arbitrary lensing applications.
-Here, we demonstrate hBN sub-wavelength resonators coupled to gold nanoantennas, compatible with standard lithographic protocols and capable of launching focused and spatially confined long-lived PhPs.
-We fabricate sub-wavelength hBN cavities at the edges between Au and Si++/SiO2 (285 nm) substrates.
-The Au edge serves as a nanoantenna, enabling coupling to the far-field component of light, while the Au/SiO2 interface acts as a partially transparent wall for PhPs: on the SiO2 side, it localizes a resonant mode within the cavity, while on the Au side, it promotes the launch of resonantly enhanced PhPs into the hBN bulk.
-The localized modes exhibit pronounced field enhancement and generate polaritons spatially defined by the cavity dimensions.
+Currently, antennas are widely used to couple the far-field component of light with polaritonic systems, facilitating the selective launch of phonon-polaritons (PhPs)21, 22, 23, 24.
+However, this approach suffers from two critical limitations: the lack of control over the wavefront and low in-plane spatial confinement of the launched PhPs.
+While hBN nanocavities exhibit remarkable mid-IR light confinement down to $\lambda$/500, where $\lambda$ is the incident wavelength, an extraordinary Purcell factor (up to 1012), and a high quality factor ($Q>400$), the resonant modes in these cavities remain trapped within the resonators25, 26, 27, 28.
+Up to our best knowledge, there has been no demonstration of utilizing these localized modes to focus propagating PhPs, although PhP focusing along the optical axis has been achieved in biaxial hyperbolic materials such as MoO3 29, where the focusing is fundamentally constrained by crystallographic directions, making it unsuitable for arbitrary-lens applications.
+This poses a significant limitation on the advancement of PhP-based technologies, and the effective control over the spatially-confined propagation and focusing in uniaxial hyperbolic materials remains elusive.
+Here, we demonstrate focused, spatially-confined, and long-lived PhPs that are launched by carefully engineered subwavelenght cavities, fabricated using standard lithographic protocols.
+We fabricate sub-wavelength hBN cavities at the edges between Au and Si++/SiO2 (285 nm) substrates (see Fig.
+1a).
+The Au edge acts as a scatterer, compensating for the momentum mismatch between incident light and PhPs.
+The Au/SiO2 interface enables the coupling of PhPs with the out-of-plane far-field component of the incident light and serves as an in-plane partially transparent mirror for PhPs.
+On the SiO2 side, it localizes a mode within the cavity, while on the Au side, it facilitates the launch of propagating and focused PhPs into the hBN bulk.
 Using scattering-type scanning near-field optical microscopy (s-SNOM), we visualize the real-space optical contrast and distinguish between cavity-launched ($\lambda_{p}$) and tip-launched ($\lambda_{p}$/2) modes.
-The former results from coupling with the far-field component of mid-IR light, which makes a significant contribution to the PhPs launched in this system.
-The far-field component, resonantly enhanced in the nano-cavities, acquires sufficient momentum to launch polaritons into the bulk of the hBN/Au heterostructures via scattering processes at the SiO2/Au interface.
-This approach contrasts with previous demonstrations of focused PhPs, where polaritons in hBN were primarily excited by the s-SNOM tip 30.
-The shapes of our resonators determine the focal points and wavefronts of the PhPs.
-By fabricating nano-resonators in various shapes, we demonstrate "concave lens"-like and "convex lens"-like focusing.
-"Concave lens"-like cavities achieve confinement down to $\lambda/55$ at the focal point.
-Thus, we propose a novel method for launching and focusing spatially confined in-plane PhPs, fully compatible with standard lithographic procedures.
+The cavity-launched PhPs result from coupling of the incident mid-IR light to the cavity modes and exhibit significant shape-dependent focusing.
+This approach for launching PhPs contrasts with s-SNOM tip-mediated excitation, which is employed in previous demonstrations of focused PhPs 30.
+We propose a new method for launching and focusing spatially confined in-plane PhPs, fully compatible with standard lithographic procedures.
+By fabricating cavities with various shapes, that determine the focal points and wavefronts of the PhPs, we demonstrate "concave lens"-like and "convex lens"-like focusing.
+Concave lens"-like cavities achieve confinement down to $\lambda/70$ at the focal point, setting a record-high confinement compared to previous studies on in-plane-focused PhPs31, 29, 32, 33.
 Results.
-In this work, we study two types of samples: (i) mechanically cleaved hBN crystals (of 30-45 nm thickness) transferred on the pre-patterned Au on Si++/SiO2 (285 nm) substrates as is (see Methods) and (ii) nanoresonators fabricated at the edges between SiO2 and Au using electron-beam lithography and plasma etching (see the schematics in Fig.1).
+In this work, we study two types of samples: (i) mechanically cleaved hBN crystals (of 30-45 nm thickness) transferred on the pre-patterned Au on Si++/SiO2 (285 nm) substrates as is (see Methods) and (ii) cavities fabricated at the interfaces between SiO2 and Au using electron-beam lithography and plasma etching (see the schematics in Fig.1).
 Fig.
@@
 A tunable ($\lambda$ = 5.5-11.5 $\mu$m) mid-infrared quantum cascade laser allows us to cover Type II hyperbolic transition - 1360-1600 cm-1 ($\epsilon_{xx}=\epsilon_{yy}<0;\epsilon_{zz}>0$) (see Methods section for details).
-Figure 2: Focusing of PhPs using patterned nanoresonantors.
+Figure 2: Focusing of PhPs using patterned subwavelength cavities.
 a) AFM topography of the Au edge with the resonators of triangular shapes.
@@
 The second wave extends from the edge of the Au stripe to the right across the entire hBN bulk (green dotted line).
-The waves, which are localized close to the edge and propagate perpendicularly to it, are the tip-launched PhPs (we do not observe hBN edge-launched fringes in our geometry)31.
+The waves, which are localized close to the edge and propagate perpendicularly to it, are the tip-launched PhPs (we do not observe hBN edge-launched fringes in our geometry)34.
 These waves originate from the mid-IR light scattered on the AFM tip.
-The observed fringes represent only a half-wavelength of the corresponding PhP ($\lambda_{PhP}$/2) since the detection moves along with the excitation 32, 33.
+The observed fringes represent only a half-wavelength of the corresponding PhP ($\lambda_{PhP}$/2) since the detection moves along with the excitation 35, 36.
 Based on this, we attribute the orange and the blue PhP waves to the tip-launched modes on Au and SiO2, respectively.
@@
 Intriguingly, they also demonstrate notably longer propagation lengths.
-We attribute these PhPs to the nanocavities-launched waves 33.
-These waves originate from the light scattered at the edge of the Au stripe.
+We attribute these PhPs to the cavity-launched waves 36.
 In the case of the green curves, the infrared light launches the corresponding PhP in the hBN on SiO2.
@@
 We find that the electric field is significantly enhanced inside the cavity exhibiting a two-fold larger reflectivity compared to the bare gold signal (and seven-fold larger compared to hBN/SiO2 signal).
-This observation establishes our nanocavities as effective PhP resonators.
+This observation establishes our subwavelength cavities as effective PhP resonators.
 To investigate PhP waves’ dependence on their shapes and sizes, we utilize a standard electron beam lithography and plasma etching (see Methods).
-We fabricate nanocavities of the desired shapes at the edges of gold patterns.
+We fabricate cavities of the desired shapes at the edges of gold patterns.
 Fig.
@@
 Attributed to an exceptional alignment (within $\sim$10 nm) between the e-beam lithography steps, we create the matching edges between the patterned hBN and the Au substrates.
-This allows us to avoid unwanted leakages of the polaritonic modes away from our nanocavities.
+This allows us to avoid unwanted leakages of the PhP modes away from our cavities.
 On the left hand side of the orange dashed line, a 45 nm thick hBN flake covers a 50 nm thick Au stripe.
@@
 Respectively, Fig.
-2b shows an experimental s-SNOM amplitude (S3) taken at $\omega$ = 1450 cm-1 (see its dispersion and the analogous set of data for a 31 nm thick hBN in Extended Data Fig.
+2b shows an experimental s-SNOM amplitude (S3) taken at $\omega$ = 1450 cm-1 (see its dispersion and the analogous set of data for a 31-nm thick hBN in Extended Data Fig.
 S3 and S4, respectively).
@@
 All calculations are done for hBN thickness of 45 nm.
-Figure 4: Concave- and convex-type PhPs focusing.
-a, b, c, d) AFM topography, numerical modeling, s-SNOM optical amplitude (S3) maps and the schematics of "concave lens"-like cavity.
-e, f, g, h) Same data for the "convex lens"-like cavity.
-Light yellow dashed lines show the edge between Au and SiO2.
-Experimental s-SNOM optical amplitude (S3) maps recorded at $\omega$ = 1410 cm-1.
+Figure 4: Convex-type PhPs focusing.
+a) AFM topography of a 2.35-$\mu$m cavity.
+The light yellow dashed line shows the edge between Au and SiO2.
 The thickness of hBN is 55 nm.
+b) Numerical modeling of the cavity.
+c) Experimental s-SNOM optical amplitude (S3) map recorded at $\omega$ = 1410 cm-1.
 Red dashed lines are a guide for eyes indicating boundaries of cavity-launched polaritons.
@@
 Green lines demonstrate tip-launched PhP waves.
+d) Schematic of the "convex lens"-like cavity operation.
+e) Experimental s-SNOM optical amplitude (S4) map of three small-diameter cavities (0.55 $\mu$m, 0.9 $\mu$m, and 1.1 $\mu$m, respectively) recorded at $\omega$ = 1405 cm-1.
+f) Dependence of the focal point distance on cavity size at $\omega$ = 1410 cm-1, as well as the confinement factor for each point.
+g) Dependence of the distance to the focal point on the excitation wavelength for small-diameter cavities (Fig.
+4e).
+The depth of focus (DoF) was determined for each cavity based on the distance oscillations.
+h) Dependence of a full width at half maximum (FWHM) (red lines) and confinement factor (green lines) on the excitation wavelength for small-diameter cavities.
 Similar to the case in Fig.
-1, two different PhP waves can be observed.
-The shorter one is a tip-launched PhP (indicated by a green arrow with the wavelength $\lambda_{PhP}/2$).
-This wave results from mid-infrared light scattered by the AFM tip, which excites phonon-polaritons that propagate toward the edge of the sample, reflect back, and interfere with themselves32, 34, 35, 36, 37.
-Interestingly, we observe the second PhP wave originating from the cavities themselves.
-Here, the observe that the peak-to-peak distance corresponds to the whole $\lambda_{PhP}$.
-It is worth noting that $\lambda_{PhP}/2$ fringes are completely suppressed in front of the cavities, and only $\lambda_{PhP}$ fringes are observed.
+1, two different PhP types can be observed.
+The one that exhibits a $\lambda_{PhP}/2$ distance between intensity maxima is a tip-launched PhP (indicated by a green arrow).
+This wave results from mid-IR light scattered by the s-SNOM tip, which excites PhPs propagating toward the edge of the sample, reflecting back, and interfering with themselves35, 37, 38, 39, 40.
+The cavity-launched PhPs show the distance between the intensity maxima of $\lambda_{PhP}$, which contrasts them with the tip-launched PhPs.
+Remarkably, $\lambda_{PhP}/2$ fringes are completely suppressed in front of the cavities, and only $\lambda_{PhP}$ fringes are observed, which implies a suppressed back-reflection of the tip-launched PhPs at the cavity interface.
 Importantly, we find that the cavity-launched PhPs acquire well-defined focal points.
-For the top cavity (cavity 1), the PhP shrinks before and expands after the focal point (marked by the red arrow).
+For cavity 1 (see Fig.
+2a,b), the PhP shrinks before and expands after the focal point (marked by the red arrow).
 The lateral size of the focal point corresponds to a confinement of $\lambda/25$.
-For the bottom two cavities, PhPs shrink with the distance away from the cavities towards their focal points that are located outside of the detectable regions.
+For cavity 2 and cavity 3, PhPs a focused at focal points outside of the detectable regions.
 The orange dashed lines serve as guides to the eye, indicating the boundaries of the cavity-launched phonon-polaritons.
 At the focal point, we identify a "hot spot" that corresponds to the electric field node.
-Consequently, the PhP’s width and the focal point location can be varied by the size of a cavity.
+This shows that the PhP’s width and the focal point location can be effectively tuned by changing the size of a cavity.
 Fig.
@@
 2b (solid lines) and their exponential decay fits (dashed lines).
-Tip-launched PhPs show a strong first peak of oscillation but a fast decay of the PhP amplitude due to high losses 38.
-On the contrary, cavity-launched PhPs exhibit a stable amplitude of oscillations with the considerably slow decay.
-As a result, pronounced optical amplitude oscillations of the cavity-launched PhPs are propagating more than three times farther than the tip-launched PhPs.
+Tip-launched PhPs exhibit rapid amplitude decay because they originate from the interference between tip-launched waves and those reflected from the edge, meaning they must complete a round-trip distance before detection.
+In contrast, cavity-launched PhPs exhibit oscillations with considerably slower amplitude decay.
+As a result, the cavity-launched PhPs propagate more than three times farther than those of tip-launched PhPs.
 […]
 Numerical simulations.
-To shed light on the physics of PhPs in our cavities and to establish an approach for their engineering with the desirable characteristics, we build a theoretical model that predicts optical properties of PhPs and effectively replicates the PhP dynamics in the triangular cavities observed in the experiment.
+To shed light on the physics of PhPs in our cavities and to establish an approach to their engineering with the desirable characteristics, we build a theoretical model that predicts optical properties of PhPs and effectively describes the PhP dynamics in the triangular cavities observed in the experiment.
 Fig.
@@
 The thickness of the hBN flake is 45 nm.
-The numerically calculated dispersion appears in a good agreement with the data extracted from the experiment albeit the dispersion for the hBN/Au interface appears to be slightly off.
-We attribute it to the buckling of the hBN flakes on the Au edge and to the surface roughness of Au; both can create gaps between hBN and Au that reduce the effective refractive index at the interface.
+The numerically calculated dispersions appear to be in good agreement with the data extracted from the experiment.
+The dispersions take into account the waviness of the hBN flake as a 5 nm gap formed due to the hBN transfer technique.
 Fig.
-3b shows the dispersion of attenuation constant Im k and compares it to the experimental data on the decay rate $\kappa=2\times\textrm{Im k}$.
+3b shows the dispersion of the attenuation constant Im k and compares it with the decay rate $\kappa=2\times\textrm{Im k}$ obtained from the experiment.
 The numerical and experimental attenuation constants well-agree with each other between 1400 cm-1 and 1460 cm-1 but diverge at the frequencies above 1460 cm-1.
-We attribute this divergence to the absorption on PMMA (e-beam lithography resist) residues that remain on the sample surface after the fabrication process.
-In this spectral range, PMMA exhibits stronger absorption with an increase in wavenumber39, 40.
-See more data on the dispersion, refractive index and the attenuation constant for the 31- and 42-nm thick hBN in Extended Data Fig.
+We attribute this divergence to the absorption of residual polymers, specifically PMMA 950 A4 (polymethyl methacrylate, an e-beam lithography resist) and PC (polycarbonate, the top surface of a dry-transfer stamp), remaining on the sample surface after the fabrication process.
+These materials exhibit high absorption in this spectral range due to the presence of various carbon bonds  41, 42, 43, 44.
+The dispersion, refractive index and the attenuation constant for the 31- and 42-nm thick hBN can be found in Extended Data Fig.
 S5 and S6.
-We use the complex effective refractive index extracted from the numerical data in Fig.
-3a,b to construct a 2D model.
-This model captures the PhP wave dynamics in the cavities and waveguides formed at the hBN/SiO2 and hBN/Au interfaces, correspondingly.
+We use the complex effective refractive index of PhPs extracted from the dispersions in Fig.
+3a,b to construct a 2D model that captures the PhP wave dynamics in the cavities and waveguides formed at the hBN/SiO2 and hBN/Au interfaces, correspondingly.
 As can be seen from the insets in Fig.
-3b, the simulated field distributions are in good agreement with the experimental optical amplitude maps.
-However, in the experiment, the optical amplitude along the gold edge is noticeably higher, which can also be attributed to the buckling effect that cannot be accounted for in our 2D model.
+3a, the simulated field distributions showcase a reasonable agreement with the experimental optical amplitude maps.
+The discrepancies between the simulated and experimental distributions, that are observed at the edges of the cavities, are due to the interactions of the s-SNOM probe and edges, while the probe is next to the edge of the cavity 45.
+Notably, this interaction is wavelength dependent.
+However, in the experiment, the optical amplitude along the Au and hBN edges is noticeably higher, which is attributed to 3D effects at the edges of the heterostructure and buckling of the hBN at the Au edge which cannot be described within our 2D model.
 Fig.
 3c demonstrates the spatial distributions of the PhPs localized in the triangular hBN/SiO2 cavities and propagating in the hBN/Au waveguide.
-The distributions show that the triangular cavities can support localized PhP modes that "leak" into the hBN/Au environment and that the leaked PhP waves are focused at a distance from the triangle base depending on the triangle size, which is also observed in the experiment (see Fig.
+The distributions show that the triangular cavities can support localized PhP modes that leak into the hBN/Au environment and that the PhP waves are focused at a distance from the triangle base depending on the triangle size, which is also observed in the experiment (see Fig.
 2b).
-This observation indicates that the optical phenomena in phonon-polaritons excited within a 3D structure can be effectively described by a 2D model.
 The schematic analysis of the triangular and rectangular cavities (see Extended Data Fig.
@@
 Fig.
-4a, 4b, and 4c show the AFM topography, numerically calculated electric field $(\textrm{Re E}_{\textrm{z}})^{2}$, and the experimental near-field optical amplitude (S3) of a submicron "concave lens"-like cavity, respectively.
-This patterned shape exhibits defocusing of the polaritons.
-We provide an intuitive explanation for this process in Fig.
-4d.
-The PhP waves launched by the Au edge, propagate perpendicular to the Au edge inside the cavity and reflect back from the concave cavity edge.
+4a, 4b, and 4c show the AFM topography, the numerically calculated electric field $(\textrm{Re},E_{\textrm{z}})^{2}$, and the experimental near-field optical amplitude (S3) of a 2.35-$\mu$m "convex lens"-like cavity, respectively.
+This patterned shape exhibits focused PhPs at a point beyond the cavity, similar to triangular cavities.
+A focal point can be observed where the curvature of the wavefront changes from positive to negative.
+Fig.
+4d illustrates a schematic of the "convex lens"-like focusing.
+The PhP waves launched by the Au edge propagate perpendicular to the Au edge inside the cavity and reflect back from the convex cavity edge.
 Upon reaching the cavity’s edge, they reflect normally to the curvature of the reflecting plane.
-The negative curvature of the reflecting plane results in the negative curvature of the polariton wavefront and, thus, defocusing.
-Figs.
-4e-g show a cavity that operates in a "convex lens" regime, focusing polaritons at a point distant from the cavity, similar to the triangular cavities.
-One can observe a focal point where the curvature of the wavefront changes sign from positive to negative as it passes through the focal point.
+The positive curvature of the reflecting plane results in a positive curvature of the PhP wavefront.
 Fig.
-4h illustrates a schematics of the "convex lens"-like focusing.
-The position of the focal point and the confinement may be adjusted by the size of the cavity.
-For a 0.6-$\mu$m-radius cavity, the confinement reaches $\lambda/55$ (see Extended Data Fig.
+4e shows the s-SNOM optical amplitude (S4) map of three cavities with different diameters: 0.55 $\mu$m, 0.9 $\mu$m, and 1.1 $\mu$m.
+The focal points are located at varying distances from the cavity/Au interfaces which is determined by its diameter.
+The focal distance as a function of the cavity diameter for $\omega=1410$ cm-1 is shown in Fig.
+4f.
+As can be seen, the focal distance linearly depends on the cavity diameter (i.e., on the curvature of the reflective plane).
+Fig.
+4g demonstrates the distance to the focal point as a function of the frequency of the incident light for three cavities illustrated in Fig.
+4e.
+The focal distance oscillates with a period specific to each cavity.
+Based on this, we determine the depth of field (DoF) for these cavities.
+DoF for classical lenses is the range of distances over which an object remains in focus, given by $DoF=n\lambda/{(NA)^{2}}$, where $n$ is the refractive index, $\lambda$ is the wavelength, and $NA$ is the numerical aperture.
+Each PhP lens has a specific focal length that depends on the geometry of the cavities, but at the same time, DoF for these lenses is finite and depends on the geometry of the cavities as well.
+By changing the excitation wavelength, we change the real-space PhP field distribution.
+Therefore, for some wavelengths, the maximum of the PhP field is located exactly in the middle of the focal depth, while for other wavelengths, it is shifted to the edge of field (see the detailed investigation in Extended Data Fig.
 S8).
-Furthermore, several cavities launching criss-crossing PhPs can be used for the formation of the interference patterns (see Extended Data Fig.
-S9) forming a lattice of $(\textrm{Re E}_{\textrm{z}})^{2}$ minima and maxima with a period equal to the polariton wavelength at the intersection.
-Thus, using the interference of cavity-launched polaritons with various wavefronts, it is possible to form a variety of polaritonic lattices for structured light-matter interaction and opto-mechanical purposes.
+Fig.
+4h demonstrates the full width at half maximum (FWHM) of the focal point as well as the confinement factor calculated as $\lambda/\textrm{FWHM}$ for cavities 1, 2, and 3 at different wavelengths.
+The FWHM decreases and the confinement increases with the increasing frequency, which is in agreement with the effective refractive index of hBN that grows with frequency (see Extended Data Fig.
+S6).
+A confinement factor as high as 70 can be achieved for the 0.9-$\mu$m cavity at $\omega$ = 1420 cm-1, placing it among the highest reported values.
+Using the same approach, we developed a "concave lens"-like cavity (see Extended Data Fig.
+S9).
+The negative curvature of the reflecting plane produces a divergent PhP wavefront and defocusing.
+Furthermore, multiple cavities launching crisscrossing PhPs can be used for the formation of interference patterns (see Extended Data Fig.
+S10), forming a lattice of $(\textrm{Re},E_{\textrm{z}})^{2}$ minima and maxima with a period equal to the PhP wavelength at the intersection.
+Thus, using the interference of cavity-launched PhP with adjusted wavefronts, it is possible to engineer polaritonic lattices for structured light-matter interaction.
 Discussion.
-To accelerate the technological application of spatially defined PhP modes, it is necessary to find a way to precisely control their spatial distribution and focus them at specific real-space locations.
-Some substantial progress has been made in tuning and focusing of the PhPs using in-plane anisotropic hyperbolic materials such as MoO341, 42, 43.
-In the case of the uniaxial hyberbolic hBN crystals, most efforts have involved a use of substrates with an in-plane anisotropy in the relevant spectral ranges 44, 45, 46.
-However, the influence of these substrates is generally limited.
-An elegant solution has been proposed by Chaudhary et.
-al.
-30 by using phase-fluid materials (Ge3Sb2Te6 on a CaF2 substrate) that locally change their dielectric constants and thus facilitate polaritonic waveguides and metalenses.
-However, the focusing effect remained rather hindered and the scalable application is somewhat elusive.
-In conclusion, we have developed hBN sub-wavelength resonators that are compatible with standard lithographic techniques, enabling the launch of focused, long-lived PhPs.
-We fabricate the subwavelength hBN cavities on the interfaces between Au and SiO2 partially transparent for PhPs.
-The localized modes showed a significant increase of the electric field and launched polaritons that remain confined in the real-space defined by the shapes and sizes of the cavities.
+To accelerate the technological application of spatially defined PhP modes, it is crucial to develop a method for precisely controlling their spatial distribution and focusing them at specific real-space locations.
+Significant progress has been made in tuning and focusing PhPs using in-plane-anisotropic hyperbolic materials such as MoO3 46, 47, 31.
+In the case of uniaxial hyperbolic hBN crystals, most efforts have relied on substrates with in-plane anisotropy in the relevant spectral ranges 48, 49, 50.
+A particularly elegant solution was proposed by Chaudhary et al.
+30, employing a phase-change material (Ge3Sb2Te6 on a CaF2 substrate) to locally modify dielectric constants and focus tip-launched PhPs.
+However, the influence of such substrates is generally limited, and integrating such uncommon materials into standard silicon-based technology remains challenging.
+Consequently, the focusing effect remains constrained, and scalable applications are elusive.
+One of the most widely adopted methods for focusing PhPs and coupling them to the far-field component of light is the use of metallic scatterers, typically made of gold 51.
+This approach is essential because the tip must not be present in operational devices.
+By employing scatterers with curved surfaces, PhPs can be focused via wave interference.
+However, this approach has a fundamental limitation: the strength of light coupling to PhP modes depends heavily on the angle between light polarization and the normal to the scatterer surface at the launch point.
+Consequently, curved scatterers can only efficiently focus PhP modes within a narrow angular range 51, making them unsuitable for achieving arbitrary PhP lensing in isotropic materials like hBN.
+Our approach overcomes these limitations by utilizing standard, technologically accessible materials (Si/SiO2, Ti/Au) as a platform for PhP engineering.
+We introduce hBN subwavelength cavities at the Au/SiO2 interface, which enable the generation of strongly focused (down to $\lambda/70$) and spatially confined PhPs.
+Crucially, our approach resolves the "scatterer" issue mentioned earlier: the Au/SiO2 interface facilitates coupling to far-field light but does not actively participate in the focusing process itself.
+Instead, the focusing mechanism is governed by the selection of PhPs with specific $k$-vector distributions inside the cavity.
+As a result, for a fixed Au/SiO2 interface orientation, cavities of varying shapes and sizes can produce different focusing effects (e.g., convex, concave, or straight wavefronts) and different angular ranges of collection.
+Thus, we believe our approach addresses several persistent challenges in PhP photonics and paves the way for numerous future discoveries.
+In conclusion, we have developed hBN sub-wavelength cavities that are compatible with standard lithographic techniques, enabling the launch of focused, long-lived PhPs.
+We fabricate the subwavelength hBN cavities at the interfaces between Au and SiO2 partially transparent for PhPs.
+The localized modes exhibit a significant enhancement of the electric field inside the cavities, launching PhPs that remain spatially confined, with their confinement determined by the shape and size of the cavities.
 Using s-SNOM, we studied the real-space optical contrast of the cavity-launched PhPs.
 Based on the PhP’s wavelength, we demonstrated coupling to the far-field component of the mid-IR light.
-The far-field component, which is resonantly enhanced in the cavities, gains enough momentum to launch polaritons into the bulk of the hBN/Au heterostructures through the scattering process at the SiO2/Au interfaces.
-The shapes of our resonators determine the focal point location as well as the wavefront shapes.
-By creating different forms of nano-resonators, we showed a focusing effect similar to a "concave lens" and a "convex lens." In particular, the "concave lens"-like cavities enable field confinement down to $\lambda/55$ at the focal points.
-By this, we realize a novel technique for generating and focusing spatially confined easy-plane PhPs compatible with the standard silicon technology.
+The far-field component, which is resonantly enhanced in the cavities, acquires sufficient momentum to launch polaritons into the bulk of the hBN/Au heterostructures through the scattering process at the SiO2/Au interfaces.
+The shape of our resonators determines both the focal point location and the wavefront geometry.
+By creating different forms of cavities, we showed a focusing effect similar to a "concave lens" and a "convex lens." In particular, the "convex lens"-like cavities enable field confinement down to $\lambda/70$ at the focal point.
+By this, we realize a novel technique for generating and focusing spatially confined in-plane PhPs compatible with the standard fabrication approaches.
 Methods
@@
 hBN flakes were prepared via mechanical exfoliation using PDMS adhesive tapes and were transferred on Si/SiO2 substrates for selection.
-Relying on optical contrast, we selected flakes of suitable thicknesses and transferred them on the pre-patterned Si++/SiO2 substrates using the PMDS/PC dry transfer method47, 48.
+Relying on optical contrast, we selected flakes of suitable thicknesses and transferred them on the pre-patterned Si++/SiO2 substrates using the PMDS/PC dry transfer method52, 53.
 For the cavity fabrication, we used the Raith EBPG5200 electron beam lithographer and PMMA 950 A4 e-beam resist.
@@
 Near-field optical measurements.
-We used a commercially available scattering-type scanning near-field microscope (s-SNOM) developed by Neaspec/Attocube (cryo-neaSCOPE) with a pseudo-heterodyne mode to a detect a background-free optical response ($\Omega\pm NM$) using the decoupled optical amplitude and phase49, 50, 51.
+We used a commercially available scattering-type scanning near-field microscope (s-SNOM) developed by Neaspec/Attocube (cryo-neaSCOPE) with a pseudo-heterodyne mode to a detect a background-free optical response ($\Omega\pm NM$) using the decoupled optical amplitude and phase54, 55, 56.
 PtIr-coated AFM tips (Nanoworld, 23 nm coating) with a resonant frequency $\approx$ 250 kHz were used to obtain the topography and the optical maps.
@@
 To describe the optical phenomena in PhPs at the hBN/SiO2 and hBN/Au interfaces, we employ numerical calculations in COMSOL Multiphysics, Wave Optics module.
-The dispersions of the propagation and attenuation constants of the PhP modes alongside the complex effective refractive indices of these modes are calculated in a 3-layer structure, where the top layer is presented by air, the second layer is presented by hBN which is an anisotropic (along the hard axis) material with permittivity tensor taken from 52 and optical axis directed perpendicular to the layers, and the third layer is either SiO2 or Au with refractive indices taken from 53, 54.
+The dispersions of the propagation and attenuation constants of the PhP modes alongside the complex effective refractive indices of these modes are calculated in a 4-layer structure, where the top layer is presented by air, the second layer is presented by hBN which is an anisotropic (along the hard axis) material with permittivity tensor taken from 57 and optical axis directed perpendicular to the layers, the third layer is a narrow air gap, and the fourth layer is either SiO2 or Au with refractive indices taken from 58, 59.
 The obtained effective refractive indices are used as refractive indices for hBN/SiO2 and hBN/Au domains in the 2D model, and the surrounding space is treated as a dielectric with a refractive index of 1.